OpenStudy (itsbribro):

The value of y varies directly with x^2, and y = 36 when x = 3. What is the value of y when x = 5?

OpenStudy (itsbribro):

@stonewoods can you please help :)

OpenStudy (itsbribro):

@rational

OpenStudy (anonymous):

Ok, so x^2=y When y=36, x=6 So when x=3 y=x^2 3^2=y y=9

OpenStudy (anonymous):

when x=5 Y=5^2 Y=25

OpenStudy (itsbribro):

OMG THANK YOU SO MUCH @Mrs.Allison <3 :) you are a life saver

OpenStudy (anonymous):

Don't worry,here to help, that's what teachers do

OpenStudy (mrnood):

@mrs.allison This is another incorrect answer @itsbribro be careful of accepting answers (in particular fromMrs.allison - she has many wrong answers to her name) y= kx^2 (this is the meaning of varies directly with x^2) when x= 3 y = 36 SO 36 = k (3^2) = k*9 so k=4 when x=5 y= 4 (5^2) = 4*25 = 100